OpenStudy (anonymous):

What is the equation of an ellipse with foci at (-4, 0) and (-4, -6) and major axis of 10? I believe my answer is wrong so please help me :)

OpenStudy (anonymous):

please help ! :)

OpenStudy (amistre64):

since the x values there are the same; we now it streches along a y axis parallel line

OpenStudy (amistre64):

half the distance between foci to find the distance to the center.. 0 - -6 = 6 6/2 = 3 the center is at (-4,-3) then

OpenStudy (amistre64):

i believe "major axis of 10" means the distance from tip to tip; so that leaves us with 5 from center to the outer vertex

OpenStudy (amistre64):

x^2 y^2 ---- + ---- = 1 s^2 L^2 L^2 = 5^2 = 25 x^2 y^2 ---- + ---- = 1 s^2 25 s^2 = Maj axis distance^2 - foci distance^2 s^2 = 5^2 - 3^2 s^2 = 25 - 9 = 16 ---------------------- x^2 y^2 ---- + ---- = 1 16 25 now when x = 0; y = +-5, and when y=0 x=+-4; but we have to put back over in its original spot; right now its at the origin

OpenStudy (amistre64):

(x+4)^2 (y+3)^2 -------- + -------- = 1 16 25

OpenStudy (amistre64):

the 16 and 25 are backwards :)

OpenStudy (amistre64):

no they arent...y is the major; never mind that idiot in the back of my head :)

OpenStudy (dumbcow):

Here is a ref sheet for ellipses, hyperbolas that might help you

OpenStudy (amistre64):

wha!?! theres already stuff out for this ;)

OpenStudy (anonymous):

Thank you both so much! That all really helped me!! I appreciate all of the help :)